What will be the acceleration of a 50 kg sledge if the tension in the rope pulling it is 200 N and frictional force is 40 N?


1
Expert's answer
2020-01-08T09:40:46-0500
"a=\frac{F}{m}=\frac{T-F_f}{m}"

"a=\frac{F}{m}=\frac{200-40}{50}=3.2\frac{m}{s^2}"